OpenStudy (anonymous):

Joan has $3.50 in nickels, dimes and quarters. She has the same amount of nickels as quarters. She has twice as many nickels as dimes. Part A. Write an equation that relates the number of each kind of coin to the amount of money that Joan has, show your work

OpenStudy (anonymous):

Please help!

OpenStudy (anonymous):

14 quarters, 70 nickels, and 35 dimes. I jsut dont know how to make the equation!!!

OpenStudy (anonymous):

Huh? If you plug in your answer the total amount is: 14*25 + 70*5 + 35*10 = $10.50 with the data you gave us the answer is: Standard form: 10d+5n+25q=350; n-q=0; -2d+n=0; substitute/eliminate d = -1/2n-5/2q+35 substitute/eliminate n = +1q-0 substitute/eliminate q = +10 Solution: d = 5; n = 10; q = 10; (d, n, q) = (5, 10, 10)